@WorkTheSpace Been very frustrating the last few years, too many things are hardcoded. I was hoping they'd rework the codebase when moving to Unity but apparently not.
@goyemeat@SickanDaDank@Verliswolf@G0ffThew Do you not understand what an analogy is? The calculation I gave you earlier already used 1/8192, the chance of not flipping a 1/8192 on each try is 8191/8192, and since every tries are independent of each other you multiply 8191/8192 by itself 6000 times
@The_douche11@Verliswolf@G0ffThew No one said it's dependent on the last encounter, since each encounter are independent, the chance of not getting a shiny in 6000 tries is (8191/8192)^6000 = 48.1%, so getting at least one shiny is 100-48.1 = 51.9%
@yoboielmo6600@Verliswolf@G0ffThew It would be a coin flip for each shiny hunter, yes it could be far higher or lower but that's just like how you can flip heads 10 times in a row
@goyemeat@SickanDaDank@Verliswolf@G0ffThew Not getting a shiny is rolling a 8191/8192, not getting a shiny in 6000 tries is just (8191/8192)^6000 = 48.1%, so getting at least one shiny in 6000 tries is 100-48.1 = 51.9%
@yoboielmo6600@Verliswolf@G0ffThew It doesn't add on top of each other, the probability of not getting a single shiny in 6000 tries is (8195/8196)^6000 = 48.1%, so getting at least one shiny is just 100-48.1 = 51.9%